OpenStudy (anonymous):

Solve for X: 3*4^x+1 -2 =19

OpenStudy (anonymous):

What is the exact question 3*4^(x+1) -2 =19 (Note the brackets)! If yes, then x= (ln(7/4))/ln(4)

OpenStudy (anonymous):

how did you get that ?

OpenStudy (anonymous):

& what exactly is ln ?

OpenStudy (anonymous):

it is log to the base e. Here's how. 3*4(x+1)-2=19 or, 3*4^(x+1)= 21 or 4^(x+1)=7 (dividing by 3 in both sides) Now taking log to base e (or ln) in both sides: (x+1)ln4=ln7 x+1=ln7/ln4 x=(ln7/ln4)-1 Or, x= (ln7-ln4)/ln4 x= ln(7/4)/ln4

OpenStudy (anonymous):

so ln is equal to log ?

OpenStudy (anonymous):

Well, log can have anything as it's base. In ln the base is constant e. But the properties of ln and log are same.